### v1

@@ 4.2 SPIRou radial velocities
We find that the properties of the RV GP trained on are different to those of the GP based on the RV data alone. The period of both GPs is the same, as we imposed in the priors the rotational period of the star. The amplitude of both GP is different 21 m s-1 for the GP based on the SOPHIE RVs alone, and 123 m s-1 RV GP trained on . The decay time of both GPs is significantly different, from 602 days from the GP based on the SOPHIE RVs alone, to 608 days from the RV GP trained on . These differences in the GP will produce different amplitudes of the Keplerian signal recovered. In Sect. 6, we will discuss in detail the properties of the Keplerian signal retrieved in the SPIRou data. We will show in Section 6, that recovered in the SOPHIE-RV data-set corrected with the GP using as proxy of activity it is almost the same that the derived from the SPIRou wPCA and Wapiti data-sets. Our analysis suggests that a GP trained on as proxy of activity provides indeed a good correction for the stellar activity RV jitter in the optical.
| Model | RMS | BIC | AICc | AICs | |
@@
| LBL km s-1 | wPCA | Wapiti | |
| (baseline solution) | |||
| Orbital parameters | |||
| [days] | |||
| [m s-1] | |||
| [JD] | |||
| Planet parameters | |||
| [M⊕] | |||
| [au] | |||
| Other parameters | |||
| [m s-1] |
6 Gl 410b physical properties
6.1 Gl 410b physical properties from the SPIRou RVs
We derived the Gl 410b planet candidate properties using RadVel. We calculated RadVel models for each of SPIRou RV data-sets: LBL km s-1, wPCA and Wapiti. Models with circular and eccentric orbits were considered. In all models, we started performing a maximum-likelihood fit to define the prior centers of the fit manually, then the MCMC model exploration was performed. Uniform bounds priors were utilized for: the planet’s period (between 0 and 20 days with initial guess at 6 days); the planet’s semi-amplitude (between 0 and 20 km s-1 with initial guess 6 m s-1); the SPIRou uncorrelated noise (between 0 and 10 m s-1 with initial guess 1 m s-1; and for (2459648.010 days and initial guess 2459648.0). The slope, dvdt, and the curvature, curv, were set equal to 0. We fed RadVel with the median-subtracted RV points, thus, the initial guess for the velocity zero-point was set to zero. The eccentricity was left to vary from 0 to 0.99. We used a stellar mass of 0.550.02 M⊙. The MCMC RadVel modeling used 50 walkers, 10000 steps, 8 ensembles and a minimum autocorrelation factor of 40.
Table 9 show the results of the RadVel AIC tests. In the km s-1 data the circular orbit is ”somewhat disfavored” with respect to the eccentric orbit. In the wPCA data the eccentric solution is favored, but the circular orbit is ”nearly indistinguishable”. In the Wapiti data the circular solution is favored, but the eccentric solutions is ”nearly indistinguishable”. The eccentricity distributions retrieved by RadVel are for the km s-1 data, for the wPCA data and for the Wapiti data. In all data-sets, the eccentricity is constrained to be lower than 0.4. As the circular and eccentric orbits provides statistically nearly indistinguishable solutions and the center of the eccentricity distributions is below 2 away from zero, we favor a circular orbit model. It has a lower number of free parameters and provides a good description of the data.
Table 10 summarizes the retrieved orbital and planet parameters for both the circular orbit model. Fig. 8 displays the phase-folded SPIRou RVs together with the best-fit circular orbit model for each of the SPIRou RV reductions, and the MCMC corner plots for the planet’s mass and semi-major axis. The RadVel models further confirm the detection of Gl 410b, as for all SPIRou data-sets the model without planet is “ruled-out”. Across all the SPIRou data-sets and models the period of the planet converges to a 6.02 days, thus a semi-mayor axis of the planet =0.0531 au. The wPCA and Wapiti data-sets give for the circular models similar results. The semi-amplitude retrieved in the LBL km s-1 data is larger, 5.30.7 m s-1, which translate to a of .
We note that the PCA methods filtering the LBL data are in fact decreasing the global dispersion of the all the RVs (not only those at km s-1) likely because of the diminution of the correlated noise and telluric correction residuals. However, the estimations of the planet mass with and without PCA correction methods are consistent within their 1 error bar. As the wPCA displays the second peak in the periodogram, and the time-series has the best correction of the km s-1 data, we favor the wPCA model. In the following, the wPCA circular orbital model is our baseline solution.
Taking the results altogether, SPIRou RVs analysis indicate that the candidate planet Gl 410b has a of Earth masses, orbits at a distance 0.05310.0006 astronomical units of its central star, with a period of 6.0200.004 days in a circular orbit. Gl 410b is a sub-Neptune planet.
@@ 6.1 Gl 410b physical properties from the SPIRou RVs
6.2 Gl 410b properties retrieved from the SOPHIE RVs corrected by activity.
@@
As discussed in Sect. 5.2, two methods were used to correct for the stellar activity RV jitter in the SOPHIE RV data and search for the RV signal of Gl 410b. In the first method, a RadVel Keplerian + GP model is fit to the SOPHIE RV data using the SOPHIE RV data alone. In the second method, a GP model trained using SPIRou’s longitudinal magnetic field measurements is fit to the SOPHIE RV time-series, and then a RadVel Keplerian model is fit to the RV residuals. In both methods, the Keplerian signal of Gl 410b is recovered at 6 days. The aim of this section is to check what are the planet properties derived, in particular , from the two activity correction methods and compare them to properties retrieved from SPIRou RVs.
@@
When GPs are used to correct for the stellar activity RV jitter, there is uncertainty on how good the activity correction is, and therefore, how reliable the planet’s properties derived are. Using the planet properties derived from SPIRou wPCA RVs as benchmark, we aim to explore what activity correction method provides the closer match. As the SPIRou data suggests a circular orbit, for this analysis, we only modeled circular orbits. We limited the analysis to the 2021 to 2023 SOPHIE observations, as these data have the quasi-simultaneous SPIRou measurements.
| Model | RMS | BIC | AICc | AICs | |
@@
As previously discussed, for the RadVel Keplerian + GP model using the SOPHIE RV data alone, we used an uniform distribution priors for the planet’s orbital period (0 to 20 days, i.e. we leave RadVel free to find the period), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days) and the uncorrelated noise (0.0 to 15.0 m s-1).
We tested the same broad priors on the RadVel model of the RV data corrected by activity with the GP trained on . RadVel retrieved the expected the period of 6 days. However, to have a better determination of the planet parameters, we narrowed down the uniform priors to (5 to 7 days)111111 This is justified because in the SOPHIE RV-GP (trained on Bℓ) data-set, the periodogram shows a strong peak at 6.02 days and the 6 day period is retrieved using the broad prior.. tc (2459648.03 days) and (2.0 to 20.0 m s-1).
Table 4 presented already the results of the RadVel AIC tests for the Keplerian + GP model for the GP using the SOPHIE RV data alone. Table 7 describes the results of the AIC tests for the SOPHIE RV - GP trained on the time-series. For both activity correction methods, the RadVel AIC tests “rule-out” the model without a planet. Table 12 summarizes the orbital solutions found by RadVel for both activity methods. Both methods of activity filtering suggest the presence of a sub-Neptune mass planet with period 6.02 days and semimayor axis 0.0531 au.
| SOPHIEGP | SOPHIE-GP | |
| GP on RVSOPHIE alone | GP trained on | |
@@
The SOPHIE RV RadVel GP + Keplerian model retrieves a of 7.6 m s-1 thus a of 14.5 M⊕. The SOPHIE - GP trained on time-series model retrieves a of m s-1 and of M⊕. While both methods provide a , thus a , consistent between them at the 1 level, the activity correction method that provides the closest match to the derived from the SPIRou RVs, is the correction by the GP trained on SPIRou’s . We display the model in Figure 9.
Our analysis suggests that the measurements of performed in the near-IR with SPIRou could be indeed a good activity proxy to correct for the stellar RV activity jitter in quasi-simultaneous measurements taken in the optical. The independent detection of the same signal in both SPIRou and SOPHIE data-sets gives solid support to a planet as being the responsible of the periodic variation seen in near-IR and optical RVs.
6.3 A joint model for Gl 410b using the SPIRou wPCA and the activity corrected SOPHIE data-sets
As the planet signal is recovered independently in the near-IR and in the optical, a joint model using both RV data-sets can be performed. To merge RV data from different wavelengths and instruments has risks and can be problematic, as instruments have different RV systematic noise and stellar activity is chromatic. Nevertheless, to check, we used RadVel and calculated a joint solution for the orbital parameters of Gl 410b using the wPCA SPIRou RVs and the GP activity corrected (using as activity proxy) SOPHIE RVs. We used an uniform distribution priors for the planet’s orbital period (0 to 20 days), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days) and the uncorrelated noise (0.0 to 10.0 m s-1) and (0.0 to 10.0 m s-1). We tested models with circular and eccentric orbits.
Table 9 summarises the results of the RadVel AIC tests. The model without planet is “ruled-out” by RadVel. The model with a planet in circular orbit is favoured, although a model with an eccentric orbit is “nearly indistinguishable”. We favor the circular solution because it describes correctly the data with fewer free parameters. We display the retrieved planet properties in Fig. 27 in the Appendix. The combination of the wPCA SPIRou and the activity corrected SOPHIE-GP data gives essentially the same planet properties as the wPCA SPIRou data alone. The only (small) differences are that a decrease in the period uncertainty from 0.004 to 0.003 days and a increase from M⊕ to M⊕ ( m s-1 to m s-1) in the combined data. The larger dispersion of the SOPHIEGP RVs with respect of wPCA SPIRou RVs could be responsible of the small increase on the planet’s semi-amplitude. The uncertainty in the period becomes smaller in the combined data-set because in both data-sets the planet signal is time and phase coherent.
| Model | RMS | BIC | AICc | AICs | |
@@ 6.3 A joint model for Gl 410b using the SPIRou wPCA and the activity corrected SOPHIE data-sets
7 Search for additional signals within the SPIRou wPCA data.
@@
7.1 A planet candidate at P=18.7 days ?
@@
If the 6.02 day Keplerian signal of Gl 410b (Table 10) is subtracted from the wPCA SPIRou time-series and a new GLS periodogram is calculated on the residuals, we find that the signal at 18.76 days is still present. Furthermore, its FAPboostrap level improves to the 0.3% level (Fig. 10 middle panel). This provides tentative evidence in favour for the presence of a second planet.
@@
To further test the plausibility of the existence of the signal at 18.7 days, we calculated a RadVel MCMC run using a two-planets model. For this test we set broad uniform priors for (0 to 20 days), (0 to 30 days), and and (0 and 20 m s-1). Circular and eccentric orbits models were calculated. The model priors and initial guesses are summarized in Table 17 in the Appendix. Our goal was to determine which planet solution is favored by RadVel and compare the periods blindly retrieved by RadVel with those of the periodogram analysis.
@@
The RadVel AIC model comparaison is presented in Table 15 in the Appendix. RadVel finds that is more likely the two-planet solution over the one-planet solution. In fact, the models with one planet are “ruled-out” by RadVel. RadVel favors a model which planet b has an eccentric orbit and planet c a circular orbit; and a model with circular orbits for both b and c. However, both models are “nearly indistinguishable”. We favor the model in which the two-planets are in circular orbits.
@@
In the Appendix, we provide the corner plots of the 2-planet in circular orbits model (Fig 28), a plot showing the phase-folded RVs and the best fitting 2-planet circular solution (Fig 29), and a table summarising the orbital parameters and physical properties retrieved (Table 18).
@@
RadVel indicates for the candidate signal at 18.7 days a = M⊕ ( m s-1 and a period =18.760.03 days (0.1130.001 au). This period, retrieved blindly by RadVel, corresponds to the peak in the periodogram once the RV signal of Gl 410b is removed from the data (middle panel of Fig. 10). The properties of Gl 410b are similar to those retrieved in the one-planet model, slightly decreases to 7.91.2 M⊕, which is within the 1 error of the one-planet model value. The bottom panel of Fig. 10, displays the periodogram of the residuals of SPIRou wPCA RVs after subtraction of the signal of Gl 410b and the candidate signal at 18.7 days. No peaks with significance higher than FAP=10% are left in the periodogram. We tested a RadVel model on the combined SPIRou wPCA and SOPHIEGP data-set (not shown), in this data also the two-planets in circular orbits model was favored by the AIC tests.
@@
Gl 410 has thus tantalising evidence for hosting two sub-Neptune planets in circular orbits close to 3:1 resonance. For curiosity, using the RadVel posterior distributions of the eccentric orbit models, we explored the stability of the system (Sect. G in the Appendix). We found that the system is safely stable for circular orbits and the nominal eccentricity values.
7.2 -periodogram / apodized sines analysis
@@
The results of the GLS periodogram and the 2-planet Radvel modeling are promising. To perform an independent assessment on the existence of periodicities within the SPIRou wPCA data-set, we performed a and apodize sine periodogram (ASP) analysis. The periodogram searches for several signals simultaneously, and the ASP tests the consistency in frequency, phase, and amplitude of the candidate signals. We used the periodogram technique, as defined in Hara et al. (2017). This tool is based on a sparse recovery technique called the basis pursuit algorithm (Chen et al., 1998). It aims to find a representation of the RV time series as a sum of a small number of sinusoids whose frequencies are in the input grid. It outputs a figure which has a similar aspect as a regular periodogram, but with fewer peaks due to aliasing. The peaks can be assigned a FAP, whose interpretation is close to the FAP of a regular periodogram peak.
@@
The periodogram takes several parameters in input, in particular a list of vectors, or predictor, which are fitted linearly along with the search for periodic signals. We chose for this linear base model several stellar activity indicators available within the SPIRou data: the longitudinal magnetic field () measurement, the LBL projection onto the 2nd derivative of the spectrum (d2v), the LBL projection onto the 3rd derivative of the spectrum (d3v), the LBL differential line width (dLW), and LBL the differential temperature measurement (dTEMP) with respect to a K template. The LBL activity indicators are calculated at the same time that the RVs (Artigau et al., 2022, 2024). We also added a periodic signal at the sidereal day, to account for apparent systematics at this period. The periodogram also necessitates to assume a certain form of the noise model. Following the method of Hara et al. (2020), we tryed a grid of noise models, and rank them with cross-validation and Bayesian evidence, computed with the Laplace approximation. The detailed analysis is presented in Appendix H.1.
@@
In Figure 11, we represent the periodogram of the noise model with the highest cross-validation score. We find clear evidence for the 6.02 days signal, and a FAP of . The 18.7 day signal is also present with a FAP. We note the presence of an additional 2.99 day signal with a FAP . Finally, we find a marginally significant signal at 1.3 days. The signals at 6, 18 and 3 days are consistently found by the periodograms computed: they present peaks in 100% of the top 20% of the models with highest cross validation scores.
@@
To further refine the analysis, we tested whether the candidate signals are strictly periodic, or if their frequency, phase and amplitudes show signs of variation. For this, we employed the method of Hara et al. (2022), which consists in fitting apodized sinusoidal signals to the dataset. These are wavelet-like signals, consisting of a sine multiplied by a Gaussian function, localized in time. In Appendix H.1, we show that there is good evidence that the three signals found by the periodogram are strictly periodic.
@@
In summary, besides the clear 6.02 day planet, we find in the and ASP analysis statistically significant candidates at 18.7 and 2.99 d, which are compatible with purely periodic signals. Let us note that the proximity of the 2.99 d planet to a 2:1 mean motion resonance with the 6 day planet further supports that indeed is a planet. Nonetheless, we do not label the two signals as confirmed planets, as their statistical significance is not completely sufficient.
@@
To conclude this section, we would like to insist that the detections of the signals at 18.7 and 2.99 days, although they pass several statistical tests, at the present are candidate signals because they are retrieved only on the wPCA SPIRou data. Further measurements and improvements in the methods to recover the data in the —Vtot—¡10 km s-1 region are required to fully exploit all the measurements taken and ultimately confirm or refute the planetary nature of these signals.

### v2

@@ 5.2 Method 2: SOPHIE Gaussian process modeling using SPIRou’s as an activity proxy
We find that the properties of the RV GP trained on are different from those of the GP based on the RV data alone. The period of both GPs is the same, as we imposed in the priors the rotational period of the star. The amplitude of each GP is slightly different: 21 m s-1 for the GP based on the SOPHIE RVs alone and 123 m s-1 for the GP trained on . The decay time of each GPs is significantly different, from 602 days from the GP based on the SOPHIE RVs alone to 608 days from the GP informed on . These differences in the GPs will produce different amplitudes of the recovered Keplerian signal. In Sect. 6, we discuss in detail the properties of the Keplerian signal retrieved in the SPIRou data. We show in Section 6, that the recovered in the SOPHIE data set corrected with the GP using as a proxy of activity is almost the same as the derived from the SPIRou wPCA and Wapiti data sets. Our analysis suggests that a GP informed on as proxy of activity provides a good correction for the stellar activity RV jitter in the optical.
6 Gl 410b physical properties
6.1 Gl 410b physical properties derived from the SPIRou radial velocities
We derived the Gl 410b planet candidate properties using RadVel. We calculated RadVel models for each of the SPIRou RV data sets: LBL km s-1, wPCA, and Wapiti. Models with circular and eccentric orbits were considered. In all models, we started performing a maximum-likelihood fit to define the prior centers of the fit manually and then we performed the MCMC model exploration. Uniform bounds priors were utilized for the planet’s period (between 0 and 20 days, with the initial guess at 6 days); the planet’s semi-amplitude (between 0 and 20 km s-1, with the initial guess 6 m s-1); the SPIRou uncorrelated noise (between 0 and 10 m s-1, with the initial guess 1 m s-1); and (2459648.010 days and an initial guess of 2459648.0). The slope (dvdt) and the curvature (curv) were set to zero. We fed RadVel the median-subtracted RV points, and thus, the initial guess for the velocity zero-point was set to zero. The eccentricity was left to vary from zero to 0.99. We used a stellar mass of 0.550.02 M⊙. The MCMC RadVel modeling used 50 walkers, 10000 steps, eight ensembles and a minimum autocorrelation factor of 40.
Table 15 show the results of the RadVel AICc tests. In the km s-1 data, the circular orbit is ”somewhat disfavored” with respect to the eccentric orbit. In the wPCA data, the eccentric solution is favored, but the circular orbit is ”nearly indistinguishable” from the eccentric orbit. In the Wapiti data, the circular solution is favored, but the eccentric solution is ”nearly indistinguishable” from the circular solution. The eccentricity distributions retrieved by RadVel are for the km s-1 data, for the wPCA data, and for the Wapiti data. In all data sets, the eccentricity, , is constrained to be lower than 0.4. As the circular and eccentric orbits provide solutions that statistically are nearly indistinguishable and the center of the eccentricity distributions is below 2 away from zero, we favor a circular orbit model. This model has a lower number of free parameters, and it provides a good description of the data.
Table 16 summarizes the retrieved orbital and planet parameters for the circular orbit model in each data set. Figure 8 displays the phase-folded SPIRou RVs together with the best-fit circular orbit model for each of the SPIRou RV reductions and the MCMC corner plots for the planet’s mass and semimajor axis. The RadVel models further confirm the detection of Gl 410b. As for all SPIRou data sets, the model without a planet is “ruled out”. Across all the SPIRou data sets and models the period of the planet converges to days; thus, a semi-mayor axis of the planet, , equal to 0.0531 au. The wPCA and Wapiti data sets give similar properties for the planet. The semi-amplitude retrieved in the LBL km s-1 data is larger, 5.30.7 m s-1, which translates to a of M⊕.
We note that the PCA methods filtering the LBL data do in fact decrease the global dispersion of the all the RVs (not only those at km s-1). This is likely because of the diminution of the correlated noise and telluric correction residuals. However, the estimations of the planet mass with and without PCA correction methods are consistent within their 1 error bar. As the wPCA displays the second peak in the periodogram, and the time series has the best correction of the km s-1 data, we favor the wPCA model. In the following, the wPCA circular orbital model is our baseline solution.
Taking the results altogether, the SPIRou RVs analysis indicate that the candidate planet Gl 410b has a of Earth masses, orbits at a distance of 0.05310.0006 au from of its central star, with a period of days in a circular orbit. Gl 410b is a sub-Neptune planet.
| Model | RMS | BIC | AICc | AICc | |
@@ 6.1 Gl 410b physical properties derived from the SPIRou radial velocities
| LBL km s-1 | wPCA | Wapiti | ||
| (baseline solution) | ||||
| Orbital parameters | [days] | |||
| [m s-1] | ||||
| [JD] | ||||
| Planet parameters | [M⊕] | |||
| [au] | ||||
| Other parameters | [m s-1] |
@@
6.2 Gl 410b properties retrieved from the SOPHIE radial velocities corrected by activity
@@
As discussed in Sect. 5.2, two methods were used to correct for the stellar activity RV jitter in the SOPHIE RV data and search for the RV signal of Gl 410b. In the first method, a RadVel Keplerian plus GP model was fit to the SOPHIE RV data using the SOPHIE RV data alone. In the second method, a GP model trained using SPIRou’s longitudinal magnetic field measurements was fit to the SOPHIE RV time series, and then a RadVel Keplerian model is fit to the RV residuals. In both methods, the Keplerian signal of Gl 410b is recovered at 6 days. The aim of this section is to check the derived planet properties, in particular , from the two activity-correction methods and compare them to properties retrieved from SPIRou RVs.
@@
When GPs are used to correct for the stellar activity RV jitter, there is uncertainty as to how good the activity correction is and, therefore, how reliable the derived planet’s properties are. Using the planet properties derived from SPIRou wPCA RVs as benchmark, we aim to determine which activity correction method provides the best match. As the SPIRou data suggest a circular orbit, we only used circular orbits for this analysis. We limited the analysis to the 2021 to 2023 SOPHIE observations, as these data have quasi-simultaneous SPIRou measurements.
| Model | RMS | BIC | AICc | AICc | |
@@
As previously discussed, for the RadVel Keplerian plus GP model using the SOPHIE RV data alone, we used a uniform distribution prior for the planet’s orbital period (0 to 20 days; i.e., we left RadVel free to find the period), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days), and the uncorrelated noise (0.0 to 15.0 m s-1).
We used the same broad priors on the RadVel model of the RV data corrected by activity with the GP trained on . RadVel retrieved the expected period of 6 days. However, to have a better determination of the planet parameters, we narrowed down the uniform priors to (5 to 7 days)171717This is justified because in the SOPHIE RV-GP (trained on Bℓ) data set, the periodogram shows a strong peak at 6.02 days and the 6-day period is retrieved using the broad prior., tc (2459648.03 days), and (2.0 to 20.0 m s-1).
Table 7 already presents the results of the RadVel AICc tests for the Keplerian plus GP model for the GP using the SOPHIE RV data alone. Table 10 describes the results of the AICc tests for the SOPHIE RV minus GP informed on the time series. For both activity-correction methods, the RadVel AICc tests “rule out” the model without a planet. Table 18 summarizes the orbital solutions found by RadVel for both activity methods. Both methods of activity filtering suggest the presence of a sub-Neptune mass planet with a period of 6.02 days and semimajor axis 0.0531 au.
| SOPHIEGP | SOPHIEGP | |
| GP on RVSOPHIE alone | GP informed on | |
@@
| Model | RMS | BIC | AICc | AICc | |
@@
The SOPHIE RV RadVel GP plus Keplerian model retrieves a of 7.6 m s-1, thus a of 14.5 M⊕. The SOPHIEGP informed on the time series model retrieves a of m s-1 and of M⊕. While both methods provide a , and thus a , consistent between them at the 1 level, the activity-correction method that provides the closest match to the derived from the SPIRou RVs is the correction by the GP trained on SPIRou’s . We display the model in Figure 9.
Our analysis suggests that the measurements of performed in the near-IR with SPIRou could be a good activity proxy to correct for the stellar RV activity jitter in quasi-simultaneous measurements taken in the optical. The independent detection of the same signal in both the SPIRou and SOPHIE data sets gives solid support to a planet being responsible of the periodic variation seen in near-IR and optical RVs.
6.3 A joint model for Gl 410b using the SPIRou wPCA and the activity-corrected SOPHIE data sets
As the planet signal is recovered independently in the near-IR and in the optical, a joint model using both RV data sets can be performed. Merging RV data from different wavelengths and instruments has risks and can be problematic, as instruments have different RV systematic noise, and stellar activity is chromatic. Nevertheless, to check the properties derived from a combined optical and near-IR RV data set, we used RadVel and calculated a joint solution for the orbital parameters of Gl 410b using the wPCA SPIRou RVs and the activity-corrected (using as an activity proxy) SOPHIE RVs. We used a uniform distribution priors for the planet’s orbital period (0 to 20 days), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days), the uncorrelated noise (0.0 to 10.0 m s-1), and (0.0 to 10.0 m s-1). We tested models with circular and eccentric orbits.
Table 12 summarizes the results of the RadVel AICc tests. The model without a planet was “ruled out” by RadVel. The model with a planet in circular orbit is favoured, although a model with an eccentric orbit is “nearly indistinguishable” from the model with a circular orbit. We favor the circular solution because it describes correctly the data with a model with fewer free parameters. We display the retrieved planet properties in Fig. 27 in the Appendix. The combination of the wPCA SPIRou and the activity-corrected SOPHIEGP data gives essentially the same planet properties as the wPCA SPIRou data alone. The only (small) differences are a decrease in the period uncertainty from 0.004 to 0.003 days and an increase in from M⊕ to M⊕ ( m s-1 to m s-1) in the combined data set. The larger dispersion of the SOPHIEGP RVs with respect of wPCA SPIRou RVs could be responsible of the small increase on the planet’s semi-amplitude. The uncertainty in the period becomes smaller in the combined data set because in both data sets the planet signal is time and phase coherent.
7 Search for additional signals within the SPIRou wPCA data
@@
7.1 Possibility of a planet at P=18.7 days
@@
If the 6.02-day Keplerian signal of Gl 410b (Table 16) is subtracted from the wPCA SPIRou time series and a new GLS periodogram is calculated on the residuals, we find that the signal at days is still present. Furthermore, its FAPboostrap level improves to 0.3% (Fig. 10, middle panel). This provides tentative evidence in favor for the presence of a second planet.
@@
To further test the plausibility of the existence of the signal at 18.7 days, we calculated a RadVel MCMC run using a two-planet model. For this test, we set broad uniform priors for (0 to 20 days), (0 to 30 days), and and (0 and 20 m s-1). Circular and eccentric orbit models were calculated. The model priors and initial guesses are summarized in Table 23 in the Appendix. Our goal was to determine which planet solution is favored by RadVel and to compare the periods blindly retrieved by RadVel with those of the periodogram analysis.
@@
The RadVel AICc model comparison is presented in Table 15 in the Appendix. RadVel finds the two-planet solution more likely compared to one-planet solution. In fact, the models with one planet were “ruled out” by RadVel. RadVel favors a model in which planet b has an eccentric orbit and planet c has a circular orbit and a model in which both b and c have circular orbits. However, the two models are “nearly indistinguishable” for RadVel. We favor the model in which the two planets are in circular orbits.
@@
In the Appendix, we provide the corner plots of the model with two planets in circular orbit (Fig 28), a plot showing the phase-folded RVs, and the best fitting two-planet circular solution (Fig 29). We also show a table summarising the retrieved orbital parameters and physical properties (Table 24).
@@
RadVel indicates for the candidate signal at 18.7 days a M⊕ ( m s-1) and a period days ( au). This period, retrieved blindly by RadVel, corresponds to the peak in the periodogram once the RV signal of Gl 410b is removed from the data (middle panel of Fig. 10). The properties of Gl 410b are similar to those retrieved in the one-planet model. We note that slightly decreases to 7.91.2 M⊕, which is within the 1 error of the one-planet model value. The bottom panel of Fig. 10 displays the periodogram of the residuals of SPIRou wPCA RVs after subtraction of the signal of Gl 410b and the candidate signal at 18.7 days. No peaks with a significance higher than FAP=10% are left in the periodogram. We also tested a RadVel model on the combined SPIRou wPCA and SOPHIEGP data set (not shown). In this data set, the model with two planets in circular orbit was favored by the AICc tests as well.
@@
Thus, Gl 410 has tantalising evidence for hosting two sub-Neptune planets in circular orbits close to a 3:1 resonance. For curiosity, using the RadVel posterior distributions of the eccentric orbit models, we explored the stability of such a system (Sect. 30 in the Appendix). We found that the system is safely stable for circular orbits and the nominal eccentricity values.
7.2 and apodized sine periodogram analysis
@@
The results of the GLS periodogram and the two-planet RadVel modeling are promising. To perform an independent assessment on the existence of periodicities within the SPIRou wPCA data set, we performed a and apodized sine periodogram (ASP) analysis. The periodogram searches for several signals simultaneously, and the ASP tests the consistency in frequency, phase, and amplitude of the candidate signals. We used the periodogram technique as defined in Hara et al. (2017). This tool is based on a sparse recovery technique called the basis pursuit algorithm (Chen et al., 1998). It aims to find a representation of the RV time series as a sum of a small number of sinusoids whose frequencies are in the input grid. It outputs a figure that has a similar aspect as a regular periodogram but with fewer peaks due to aliasing. The peaks can be assigned an FAP, whose interpretation is close to the FAP of a regular periodogram peak.
@@
The periodogram takes several parameters as input, in particular, a list of vectors, or predictor, that are fitted linearly along with the search for periodic signals. For this linear base model, we chose several stellar activity indicators available within the SPIRou data: the longitudinal magnetic field () measurement, the LBL projection onto the second derivative of the spectrum (d2v), the LBL projection onto the third derivative of the spectrum (d3v), the LBL differential line width (dLW), and LBL the differential temperature measurement (dTEMP) with respect to a K template. The LBL activity indicators are calculated at the same time as the RVs (Artigau et al., 2022, 2024). We also added a periodic signal at the sidereal day to account for apparent systematics at this period. The periodogram also necessitates to assume a certain form of the noise model. Following the method of Hara et al. (2020), we tested a grid of noise models, and ranked them with cross-validation and Bayesian evidence, which we computed with the Laplace approximation. The detailed analysis is presented in Appendix G.1.
@@
In Figure 11, we represent the periodogram of the noise model with the highest cross-validation score. We found clear evidence for the 6.02-day signal with an FAP of . The 18.7-day signal is also present with a FAP. We note the presence of an additional 2.99-day signal with an FAP . Finally, we found a marginally significant signal at 1.3 days. The signals at 6, 18, and 3 days are consistently found by the computed periodograms, they present peaks in 100% of the top 20% of the models with highest cross-validation scores.
@@
To further refine the analysis, we tested whether the candidate signals are strictly periodic or if their frequency, phase, and amplitudes show signs of variation. For this, we employed the method of Hara et al. (2022), which consists in fitting apodized sinusoidal signals to the data set. These signals are wavelet-like, composed of a sine multiplied by a Gaussian function localized in time. In Appendix G.1, we show that there is good evidence that the three signals found by the periodogram are strictly periodic.
@@
In summary, apart from the clear 6.02-day planet, we find in the and ASP analysis statistically significant candidates at 18.7 and 2.99 d, which are compatible with purely periodic signals. We note that the proximity of the 2.99 d planet to a 2:1 mean motion resonance with the 6-day planet further supports that it is indeed is a planet. This third planet would have a mass of a few Earth masses. Nonetheless, we do not label the two signals as confirmed planets, as their statistical significance is not completely sufficient.
@@
To conclude this section, we insist that although the detections of the signals at 18.7 and 2.99 days pass several statistical tests, at the present they are candidate signals because they are only retrieved in the wPCA SPIRou data. Further measurements and improvements on the methods to recover the data in the —Vtot—¡10 km s-1 region are required to fully exploit all the measurements taken and to ultimately confirm or refute the planetary nature of these signals.